# Non-uniqueness of the steady state for run-and-tumble particles with a double-well interaction potential

### v4

@@ C.1 Numerical study of a case where the interaction force vanishes at infinity
with $a>0$ and $n$ an even integer larger or equal to $4$, and performed numerical simulations for $N=100$ particles, see Fig. 10. The results shown here are for $n=4$, but other values of $n$ give similar results. We find that, as long as the absolute value of the minimum of the force is much larger than the value of the maximum, the shape of the density in the stationary state is qualitatively similar to what we found for $a=0$. In particular, we still observe a transition between a connected and a disconnected support when varying the parameter $k_{0}$.
@@ C.2 Limit $\gamma\to 0$
Example. One example is $W(x)=\frac{1}{2}\log(1+x^{2})$, $W^{\prime}(x)=-x/(1+x^{2})$. One finds $y=\frac{1}{8v_{0}}(1-\sqrt{1-16v_{0}^{2}}\,)$ which converges to $y\to y_{m}=1/2$ when $v_{0}\to v_{m}=1/4$. We have confirmed this in a numerical simulation with $N=100$ particles and various small values of $\gamma$. For $v_{0}>v_{m}$, we observe a transient regime, after which the particles escape to infinity (see the right panel of Fig. 11). The duration of the transient state does not seem to depend on $N$, but increases with $\gamma$. This can be understood as for smaller $\gamma$ the motion is more persistent which helps the particle to escape.

### v5

@@ C.1 Numerical study of a case where the interaction force vanishes at infinity
with $a>0$ and $n$ an even integer larger or equal to 44, and performed numerical simulations for $N=100$ particles, see Fig. 10. The results shown here are for $n=4$, but other values of $n$ give similar results. We find that, as long as the absolute value of the minimum of the force is much larger than the value of the maximum, the shape of the density in the stationary state is qualitatively similar to what we found for $a=0$. In particular, we still observe a transition between a connected and a disconnected support when varying the parameter $k_{0}$.
@@ C.2 Limit $\gamma\to 0$
Example. One example is $W(x)=\frac{1}{2}\log(1+x^{2})$, $W^{\prime}(x)=-x/(1+x^{2})$. One finds $y=\frac{1}{8v_{0}}(1-\sqrt{1-16v_{0}^{2}}\,)$ which converges to $y\to y_{m}=1/2$ when $v_{0}\to v_{m}=1/4$. We have confirmed this in a numerical simulation with $N=100$ particles and various small values of $\gamma$. For $v_{0}>v_{m}$, we observe a transient regime, after which the particles escape to infinity (see the right panel of Fig. 11). The duration of the transient state does not seem to depend on NN, but increases with $\gamma$. This can be understood as for smaller $\gamma$ the motion is more persistent which helps the particle to escape.